OpenStudy (anonymous):

xy'=y+2*sqrt(xy) Somewhere in there is a homogeneous equation yearning to be free, I can't find it though..

OpenStudy (anonymous):

are you needing the integral?

OpenStudy (anonymous):

The general solution, the ex bank i'm running though are all first order DQs. normally, i'd v=y/x, but that root throws me off.

OpenStudy (anonymous):

dont think of it as a root then. sqrt (x)= x^(1/2)

OpenStudy (anonymous):

then dividing both sides, y'=y/x + 2(xy)^(1/2)/x

OpenStudy (anonymous):

I'm attempting to format this a a homeogeneous FODQ but perhaps I don't know all the tricks of the trade

OpenStudy (anonymous):

which can be rephrased. y'=y*x^(-1)+2*y^(1/2)*x^(-1/2)

OpenStudy (anonymous):

its easier looking at all powers rather than sqrts and fractions

OpenStudy (anonymous):

so then that's 1/2 v ? for the subsitution \[v+v'x = v + 2 v^(1/2)\]

OpenStudy (anonymous):

then I can solve by seperation of variables correct?

OpenStudy (anonymous):

what are you setting as v?

OpenStudy (anonymous):

I'm subbing v=y/x so y'=v +v'x

OpenStudy (anonymous):

ok. yes that would be v^(1/2) then. you good from there?

OpenStudy (anonymous):

appears so, thank you, now i'll just see if works out in the end

OpenStudy (anonymous):

lol. good luck. simple things like looking at the same number in a different way does wonders.

OpenStudy (anonymous):

cherrio,y=xln(x)^2 +xc

OpenStudy (anonymous):

again - its hfodeq if y'=f(x,y) where f(x,xt) = f(1,t)